### v1

@@ 2.7 Summary of microbiology part of the SAP-1
Ground-based experiments have been conducted to observe the growth of Bacillus clausii. Through various trials, it has been determined that maintaining a 75% aeration volume in the culture environment results in a more significant final growth. This setup provides a reliable range for Optical Density (OD) measurements over an extended duration.
@@ 3.1 Overview
The mechanical instrumentation segment encompasses the comprehensive configuration of the SAP-1 payload, addressing structural, thermal, and logistical considerations. SAP-1 autonomously executes essential experimental procedures in a distinctive yet straightforward manner. The system features a rotating platform, characterized by axes of symmetry passing through the center and connected to a motor (refer Figure 4). Positioned at the distant ends of the rotating platform are bio-wells containing bacteria in their sporulated form. In the center of the platform, a tank holds the required nutrient broth for spore germination and growth. The tank is maintained at higher pressure, with a piston segregating pressurized air from the nutrient liquid. The bio-wells, at a relatively lower pressure compared to the tank, are connected to the tank through solenoid valves and plumbing. The entire setup is housed within a hermetically sealed chamber, employing appropriate heating and temperature control techniques to maintain Earth-like pressure and temperature conditions.
@@
Upon command, the platform initiates rotation, and valves open for a short, predetermined interval, enabling a specific amount of nutrient broth to flow into the bio-wells. Subsequently, the platform executes an agitation cycle (jittery motion) to ensure a uniform mixing of bacteria with the nutrient medium. During the growth phase, continuous stirring of the bacterial culture is crucial, necessitating similar motor actions for periodic mixing of the cultured media. The lack of gravity in LEO poses challenges in obtaining pure Optical Density (OD) readings of a liquid-air mixture floating in space. To address this, the platform is given a constant angular velocity during the OD Measurement instants, inducing centrifugal forces to force the culture at the farthest end of the bio-well. The hermetically sealed chamber, maintained at 1 bar pressure, is filled with Nitrogen gas. The rotation of the platform also aids in uniformly distributing heat throughout the chamber.
@@ 3.3 A brief insight
The complete rotating platform is encapsulated within a hermetically sealed chamber, which functions as the foundation for the rotating platform and serves as the core support structure for the entire SAP-1 payload [refer Figure 6]. The final dimensions of the entire setup are targeted to be 1U (10cm X 10cm X 10cm) with its weight being under 1500 grams.
@@ 3.3.1 Tests, simulations and analyses
Structural analysis and thermal analysis of the technology demonstration model of the SAP-1 payload were conducted using simulation software named SOLIDWORKS (by Dassault Systèmes) and Ansys Thermal Desktop respectively, with satisfactory results [refer Figure 7].
Fig. 7: Resulting image of a computer simulated displacement analysis of the rotating platform of SAP-1 (Structural analysis using SOLIDWORKS).
3.4 Summary - Mechanical instrumentation segment
The mechanical instrumentation segment of the SAP-1 payload involves a sophisticated design addressing structural, thermal, and logistical aspects. It employs a rotating platform mechanism with bio-wells containing bacteria, a nutrient tank under pressure, and a hermetically sealed chamber with temperature control for Earth-like conditions in space.
The platform’s rotation initiates experimental procedures, ensuring uniform mixing of bacteria with nutrient broth and preventing bacterial lumping. In the absence of gravity, the platform’s motion induces centrifugal forces, aiding in obtaining accurate OD measurements. The hermetically sealed chamber, pressurized with Nitrogen gas, facilitates uniform heat distribution.
@@ 4.1 Overview
The electronics instrumentation segment serves as the backbone of the entire operational framework of SAP-1 [refer Figure 8]. It plays a crucial role in commanding and controlling the payload’s subsystems, as well as facilitating the relay of information to and from ground control. The primary functionalities of the electronics segment can be enumerated as follows:
@@
    Communicating the data generated to and accepting manual commands from ground station
@@
    Acquiring, storing and redistributing electrical power as per requirements and availability
The electronics segment of SAP-1 can be classified into the following parts:
@@ 4.2 Control subsystem
Valve control system: The valves employed in SAP-1 are solenoid valves, and due to the inductive nature of solenoids, they introduce some transience to step inputs. Consequently, to instantaneously activate a solenoid valve, the magnetic flux of the solenoid’s inductor must be rapidly ’magnetically charged’ at the moment of the step input and sustained at a constant level after reaching the threshold. Achieving this requires a specialized circuit known as a ’Spike and Hold’ circuit. The fundamental function of this circuit is to supply a high input voltage, referred to as the spike voltage, to the solenoid valve for a brief duration when the command to activate the valve is received. Subsequently, it switches back to a lower voltage source, termed the hold voltage, to maintain the magnetic flux created by the spike voltage.
@@
In microbiological applications, optical density measurements are typically conducted at a wavelength of 600 nm. This specific wavelength is chosen because most substances involved in microbiological experiments, including microorganisms, exhibit minimal response or interference with light in the vicinity of 600 nm. Consequently, OD at 600 nm is a widely utilised metric in microbiological experiments [13].
@@ 4.3 Power subsystem
The power subsystem in SAP-1 serves functions akin to those found in the majority of satellites or payloads but is tailored for shorter-duration operations (around 30 Days). It is tasked with providing power to all electronic subsystems of SAP-1, which is designed to have a peak power consumption of less than 10 watts and a nominal power consumption of 5 watts. As depicted in the block diagram at the beginning of the chapter, the power subsystem receives input power from two sources: one from the power supply of the orbital experimental module and the other from solar panels. In instances of intermittent power supply from the orbital experimental module or when it fails to meet the peak power requirements of SAP-1, a battery will be incorporated to ensure a balanced and consistent power supply. Thus the power subsystem incorporates a battery management system as well.
@@ 4.5 Summary - Electronics instrumentation subsystem
The electronics segment serves as the operational backbone of SAP-1, commanding subsystems, and facilitating information exchange with ground control. It encompasses the Control, Power, and Communication subsystems.
@@
The electronics segment is vital for SAP-1’s autonomy, experimentation, data handling, and communication. The systems have been tailored for space conditions, ensuring reliable performance.
@@ 5 Conclusions
The SAP-1 initiative is dedicated to investigating the effects of microgravity and space radiation on living organisms, with a specific focus on the critical role of microbial organisms in life support. This research concentrates on the probiotic bacteria Bacillus clausii, aiming to explore its growth and survival in the space environment at LEO. The significance of this study is rooted in its potential contributions to astronaut health, considering the gut microbiota’s essential involvement in digestion, immune function, and mental health.
The importance of probiotics in space becomes apparent as we investigate their ability to alleviate the impacts of space-induced stress, gastrointestinal issues, and psychological challenges encountered by astronauts. The use of Bacillus clausii, a robust probiotic strain, emerges as a promising approach to preserving gut health, improving immune function, and potentially supporting mental well-being during prolonged space missions. The payload design within the SAP-1 initiative, incorporating a rotating platform, tackles crucial challenges in automating microbiological experiments in microgravity. This design ensures uniform mixing, prevents bacterial clumping, and facilitates density-based separation for precise optical density measurements.
The rotating platform within the mechanical instrumentation segment plays a crucial role in preserving the integrity of the experiments. It ensures even distribution of nutrient broth and uniform heat transfer within the hermetically sealed chamber. This innovative solution offers a comprehensive approach to overcoming challenges specific to microbiological experiments in space, underscoring its significance in replicating conditions akin to those on Earth.
The electronics instrumentation segment functions as the operational core, by controlling the autonomous execution of experimental procedures, managing motor and valve systems, and handling data acquisition and communication with ground control. The integrated control, power, and communication subsystems guarantee the smooth operation of SAP-1, showcasing a careful integration of technology to align with the goals of this pioneering astrobiological research.
In conclusion, the SAP-1 initiative, with its interdisciplinary approach by spanning the fields of microbiology, mechanics, and electronics, not only enhances our comprehension of probiotic behavior in space but also establishes a foundation for forthcoming missions involving microbial experiments. As we explore the uncharted realms of microgravity, the information gathered from SAP-1 is expected to have a meaningful impact on the well-being and success of astronauts in upcoming space exploration missions.

### v2

@@ 2.7 Summary of microbiology part of the SAP-1
Ground-based experiments have been conducted to observe the growth of Bacillus clausii. Through various trials, it has been determined that maintaining a 75% aeration volume in the culture environment results in a more significant final growth. This setup provides a reliable range for OD (OD) measurements over an extended duration.
@@ 3.1 Overview
The mechanical instrumentation encompasses the comprehensive configuration of the SAP-1 payload, addressing structural, thermal, and logistical considerations. SAP-1 autonomously executes essential experimental procedures in a distinctive yet straightforward manner. The system features a rotating platform, characterized by axes of symmetry passing through the center and connected to a motor (refer Figure 4). Positioned at the distant ends of the rotating platform are bio-wells containing bacteria in their sporulated form. In the center of the platform, a tank holds the required nutrient broth for spore germination and growth. The tank is maintained at higher pressure, with a piston segregating pressurized air from the nutrient liquid. The bio-wells are connected to the tank through solenoid valves and plumbing. The nutrient tank is pressurized slightly above atmospheric pressure, while the bio-wells are kept at atmospheric pressure, creating a pressure difference between the two. To ensure reliable separation of the air and nutrient broth, a piston-based system has been selected due to its ability to provide the necessary displacement force and its proven leak-proof performance under extreme conditions. This design will undergo rigorous testing for leaks to ensure reliability in space-grade applications. The entire setup is housed within a hermetically sealed chamber, employing appropriate heating and temperature control techniques to maintain Earth-like pressure and temperature conditions.
@@
Upon command, the platform initiates rotation, and valves open for a short, predetermined interval, enabling a specific amount of nutrient broth to flow into the bio-wells. Subsequently, the platform executes an agitation cycle (jittery motion) to ensure a uniform mixing of bacteria with the nutrient medium. During the growth phase, continuous stirring of the bacterial culture is crucial, necessitating similar motor actions for periodic mixing of the cultured media. The lack of gravity in LEO poses challenges in obtaining pure OD (OD) readings of a liquid-air mixture floating in space. To address this, the platform is given a constant angular velocity during the OD Measurement instants, inducing centrifugal forces to force the culture at the farthest end of the bio-well. The hermetically sealed chamber, maintained at 1 bar pressure, is filled with nitrogen gas. The rotation of the platform also aids in uniformly distributing heat throughout the chamber.
@@ 3.3 A brief insight
The complete rotating platform is encapsulated within a hermetically sealed chamber, which functions as the foundation for the rotating platform and serves as the core support structure for the entire SAP-1 payload (refer Figure 6). The final dimensions of the entire setup are targeted to be 1U (10cm X 10cm X 10cm) with its weight being under 1500 grams.
@@ 3.3.1 Tests, simulations and analyses
Structural analysis and thermal analysis of the technology demonstration model of the SAP-1 payload were conducted using simulation software named SOLIDWORKS (by Dassault Systèmes) and Ansys Thermal Desktop respectively, with satisfactory results.
3.4 Summary - Mechanical instrumentation
The mechanical instrumentation of the SAP-1 payload involves a sophisticated design addressing structural, thermal, and logistical aspects. It employs a rotating platform mechanism with bio-wells containing bacteria, a nutrient tank under pressure, and a hermetically sealed chamber with temperature control for Earth-like conditions in space.
The platform’s rotation initiates experimental procedures, ensuring uniform mixing of bacteria with nutrient broth and preventing bacterial lumping. In the absence of gravity, the platform’s motion induces centrifugal forces, aiding in obtaining accurate OD measurements. The rotating platform, with its large surface area, helps to mix the air inside the hermetically sealed chamber. This induces forced mechanical convection, allowing heat to be distributed more uniformly by promoting air movement and mixing within the chamber.
@@ 4.1 Overview
The electronics instrumentation serves as the backbone of the entire operational framework of SAP-1 (refer Figure 7). It plays a crucial role in commanding and controlling the payload’s subsystems, as well as facilitating the relay of information to and from ground control. The primary functionalities of the electronics can be enumerated as follows:
@@
    Communicating the data generated to and accepting manual commands from the ground station
@@
    Acquiring, storing, and redistributing electrical power as per requirements and availability
The electronics of SAP-1 can be classified into the following parts:
@@ 4.2 Control subsystem
Valve control system: The valves employed in SAP-1 are solenoid valves, and due to the inductive nature of solenoids, they introduce some transience to step inputs. Consequently, to instantaneously activate a solenoid valve, the magnetic flux of the solenoid’s inductor must be rapidly ’magnetically charged’ at the moment of the step input and sustained at a constant level after reaching the threshold. Achieving this requires a specialized circuit known as a ’Spike and Hold’ circuit. This is an electronic circuit that delivers a high-voltage pulse (spike) to a solenoid valve for a brief duration to activate it, followed by a lower voltage (hold) to maintain its position. This method optimizes power consumption and extends the lifespan of the solenoid valve. The spike voltage overcomes the initial resistance of the valve, while the hold voltage ensures continuous operation without excessive power dissipation. In the current prototype, the spike voltage is 5V, and the hold voltage is 3.3V.
@@
In microbiological applications, OD measurements are typically conducted at a wavelength of 600 nm. This specific wavelength is chosen because most substances involved in microbiological experiments, including microorganisms, exhibit minimal response or interference with light in the vicinity of 600 nm. Consequently, OD at 600 nm is a widely utilised metric in microbiological experiments Kallapur et al. (2021b).
@@ 4.3 Power subsystem
The power subsystem in SAP-1 serves functions akin to those found in the majority of satellites or payloads but is tailored for shorter-duration operations (around 30 Days). It is tasked with providing power to all electronic subsystems of SAP-1, which is designed to have a peak power consumption of less than 5 watts and a nominal power consumption of 5 watts. As depicted in the block diagram at the beginning of the chapter, the power subsystem receives input power from two sources: one from the power supply of the orbital experimental module and the other from solar panels. In instances of intermittent power supply from the orbital experimental module or when it fails to meet the peak power requirements of SAP-1, a battery will be incorporated to ensure a balanced and consistent power supply. Thus the power subsystem incorporates a battery management system as well.
@@ 4.5 Summary - Electronics instrumentation subsystem
The electronics serves as the operational backbone of SAP-1, commanding subsystems, and facilitating information exchange with ground control. It encompasses the Control, Power, and Communication subsystems.
@@
The electronics is vital for SAP-1’s autonomy, experimentation, data handling, and communication. The systems have been tailored for space conditions, ensuring reliable performance.
@@ 5 Conclusions
The SAP-1 initiative is dedicated to investigating the effects of microgravity and space radiation on living organisms, with a specific focus on the critical role of microbial organisms in life support. SAP-1 concentrates on the probiotic bacteria Bacillus clausii, aiming to explore its growth and survival in the space environment at LEO. The significance of this study is rooted in its potential contributions to astronaut health, considering the gut microbiota’s essential involvement in digestion, immune function, and mental health.
The importance of probiotics in space becomes apparent as we investigate their ability to alleviate the impacts of space-induced stress, gastrointestinal issues, and psychological challenges encountered by astronauts. The use of Bacillus clausii, a robust probiotic strain, emerges as a promising approach to preserving gut health, improving immune function, and potentially supporting mental well-being during prolonged space missions. The payload design within the SAP-1 initiative, incorporating a rotating platform, tackles crucial challenges in automating microbiological experiments in microgravity. This design ensures uniform mixing, prevents bacterial clumping, and facilitates density-based separation for precise OD measurements.
The rotating platform within the mechanical instrumentation plays a crucial role in preserving the integrity of the experiments. It ensures even distribution of nutrient broth and uniform heat transfer within the hermetically sealed chamber. This innovative solution offers a comprehensive approach to overcoming challenges specific to microbiological experiments in space.
The electronics instrumentation functions as the operational core, by controlling the autonomous execution of experimental procedures, managing motor and valve systems, and handling data acquisition and communication with ground control. The integrated control, power, and communication subsystems guarantee the smooth operation of SAP-1, showcasing a careful integration of technology to align with the goals of this pioneering astrobiological research.
Although the SAP-1 payload can be flown either on an independent nanosatellite or onboard a space platform, the latter option is more advantageous. This approach eliminates the need for controlling the satellite’s maneuvers, and the required power can be drawn directly from the space platform. Specifically, we plan to use SAP-1 on the Polar Satellite Launch Vehicle (PSLV) Orbital Experimental Module (POEM) [^4] More info on POEM: Webpage link.
In conclusion, the SAP-1 initiative, with its interdisciplinary approach by spanning the fields of microbiology, mechanics, and electronics, not only enhances our comprehension of probiotic behavior in space but also establishes a foundation for forthcoming missions involving microbial experiments. As we explore the uncharted research into microgravity, the information gathered from SAP-1 is expected to have a meaningful impact on the well-being and success of astronauts in upcoming space exploration missions.